An acid-base titration was performed. Calcium hydroxide was used to titrate HCl. It took only 9.16 mL of the base, Ca(OH)2, to titrate 25.00 mL of 0.985 M HCl, the acid. What was the Molarity of the Ca(OH)2?

1
Expert's answer
2020-12-06T13:55:19-0500

Ca(OH)2 + 2HCl → CaCl2 + 2H2O

"V(Ca(OH)_2) = 9.16 mL \\

C(HCl) = 0.985 \;M \\

V(HCl) = 25.0\; mL"

Proportion 1 (HCl):

0.985 – 1000 mL

n(HCl) – 25.0 mL

"n(HCl) = \frac{0.985 \times 25}{1000} = 0.024625 \;mol"

According to the reaction

n(Ca(OH)2)"= \frac{1}{2}n(HCl) = 0.0123 \; mol"

Proportion 2 (Ca(OH)2):

0.0123 mol – 9.16 mL

C – 1000 mL

C(Ca(OH)2) "= \frac{0.0123 \times 1000}{9.16} = 1.34 \;M"

Answer: 1.34 M